MODS1 Commissioning Run 5: UTC 2011 February 18

Observer(s): R. Pogge & J. Antognini (OSU), M. Pedani (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: A. Rakich & D. Miller remote from Tucson at the start
Telescope Support: J. Urban (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Full Moon tonight. Clear and cooler at sunset. Active optics system testing occupied the first 3 hours of the night. The last 6.5 hours of the night (0600-1230) were lost to high winds.

Preparations

One of our preparation goals has been to get python/pyraf/numpy working on the obsN machines so we can try out the prototype modsalign.py script for MODS MOS mask alignment. We're having the usual problems of juggling updates, so no joy yet, but Stephen Hooper is on the case so we should converge soon.

The MODS half-height computer rack is now empty and has been moved into storage inside the MODS cart in the high-bay preparatory to shipping it to OSU with the next round of stuff to go back (we need it for MODS2 lab integration and testing). We are fully in the LBT rack except for an auxiliary monitor that is on a laptop table until an appropriate rack shelf can be procured. A newer and more compact LCD monitor will be shipped out to take the place of the thing we're using now, and then we'll be fully loaded into the rack. Getting cooling going would be nice, however, as this new rack is more enclosed than the open half-height lab rack we've been using.

Instrument and Observing Software: implemented and tested some minor software tweaks from the previous night (mostly for script support and suppression of old debug messages that are no longer relevant). We are ready for tonight. Also formalized the new CCD ROI parameters for the prism MOS mode (4Kx3K readout window). We're ready to go for for tonight. The goal is to get all the offset (rotation and xy) hooks worked out, and then make what modifications to our software and the modsUI gui to implement them.

Torsten has loaded updates to the GCS and the sextractor files that should address the bright-background-as-false-guide-star problem. Andrew and Doug want to do some active optics tests at the start with IDL to check out the next set of updates to the WF analyzer so they can clear the way to getting that implemented in GCS.

Details

All times are UTC unless otherwise noted.

Afternoon

A light tree-dusting of snow last night, but now it is mostly clear with a few puffy clouds around us. Humidity is still 100% and wind is 10-15m/s, so to get anything tonight it'll have to dry out and calm down a little.

Opening up and setting Pointing

0005 - Still clouds around, but only orographic puffies. Humidity is finally dropping to around 80%, wind is an issue but still only gusting to 15m/s. The building is mostly free of snow/ice, and they're rotating the building to expose the remaining ice to the bright sun. Temp is -3°C. With the winds over 10m/s it will be challenging to get the mirrors into equilibration.

0040 - Humidity down to 70% - DGH has the side vent doors open and the mirror is starting to cool passively with a little assist from outside air. Most of the orographic clouds are gone but some junk on the horizon. It won't be cleanly photometric, but good enough to work if the wind and humidity continue their favorable trends.

0125 - Sunset (nicer tonight), much clearer above than before, humidity down in the 60's and temp -2.5°C. Wind still as before. David will be opening us up shortly and we'll get going for the night.

0142 - First Pointing Star
   IE = -60
   CA = +104
David had to dial in some focus at the start, it was a big fuzzy donut originally.

Active Optics Tests

0153 - On-axis active preset to BS9107 (sent from modsUI)

Star is out of focus, and sextractor is rejecting it, causing the preset to fail. A bit of overcompensation, eh?

David focusing by hand and running the active preset again.

This time it got to the offset onto the hotspot, and then it looks like sextractor failed to find the star. Calling Torsten... the problem is sextractor failing. GCS also hung saying "acquiring full guide CCD read out..."

0204 - Calling Dan Cox for assistance with the GCS software. Torsten is out and away from a computer right now.

0215 - Dan restored the old version - trying the preset again. Worked!

Seeing is 1.5-2 arcsec - not surprising as the mirror is still much warmer than the outside air by more than +2°C.

Going to let it run for a while until it settles down.

0226 collimating with GCS (BS9107 on axis) with bad dome seeing, convergence jumping from 200 nm - 400 nm. Going to let it run for a while, then switch to on-axis collimating with IDL.

0232 collimating on axis with idl (collimate_multiple, 20)

0233 Guide star almost jumped out of its box, Torsten, please see what guiding was doing here, it came back though.

0234 seeing ~ 2"

0236 idl appears to be diverging on-axis. Sorted, it turns out you have to write in idl "collimate_multiple, 20 /send" if you want to send corrections...

0240 collimating with idl on-axis now, then will restart WFS with GCS after some iterations of idl. Stopped at wfsc35

0300 Failed acquisition at the verify guidi8ng stage.

0301 GCS collimating. Rick tries "pause guiding" (alias gpause) and "resume guiding (gresume) which worked to pause guiding and WFS (though the WFS bar in the GCS GUI stayed green and gave no other info).

0310 GCS WFS converging to ~ 300 nm

0311 IDL collimating. First WFSC image wfsc 37 some astigmatism jump but may be dome seeing. Pupils improving but still jumpy, going back to GCS wfs.

0314 active on-axis preset with GCS, going to wait till dome seeing has really settled down.

0315 preset failed in the same place and the same way as above. As with last time, succeded on second try.

0330 Dave H takes over to rotate the building to do a wind measurement. Then active preset to BS9126 GS 0.

0345 Several good wavefronts with GCS. Switching to collimating with IDL. wfsc 42. Looks good, no big jumps in collimation. 0352 sending active preset to BS9126 off-axis gs (4.0'). Looking for a jump from on-axis.

z4 z5 z6 z7 z8
-43 353 546 293 125
55 167 288 149 141
-138 201 226 39 72
10 79 -39 -5 99
166 -108 -2 24 109
71 -14 -142 61 64
-71 -140 119 11 -24
75 86 -54 36 19

0358 collimating with IDL on same off-axis star.

sent to PSF values...

z4 z5 z6 z7 z8
-163 327 126 123 345
96 -129 218 13 134
95 7 209 13 110
-1 59 131 -17 93
-19 20 174 54 86
151 33 117 -17 -14
131 47 -78 22 54

0405 Sending guide preset on axis and collimating with IDL, hope to see no big jump...

0407 Preset failed have to redo pointing (!!!!). IE -75 CA 90.

0427 all presets between now and last log entry failed. Re-doing pointing on M5_0094, Ie -85 CA 90

0440 turns out that we had a communication error and as a result we never rolled back to yesterdays version of SExtractor parameters and GCS are there now. David rotating building to check wind.

0503 So we made a new field data list for BS9118 got started on it, then had to rotate teh building for wind to the NE. Now setting a new target list for BS9122, hopefully more wind friendly.

0519 Taking field data on BS9122 with IDL (GCS collimating on axis and taking WFSC image off axis).

0525 failed on first off-axis preset (IDLs fault this time) need some work on IDL script.

0532 resending field_collect script preset to BS9177 starting at wfsc 113

0540 - A minor glitch interrupted the script. An AGw stage "hom"e command was sent from the MODS gui in error (Rick accidentally clicked the guider home button when hitting the mouse to get the obs2 workstation out of screensaver mode). This caused a conflict between the IDL script and the modsGUI which were both hitting the AGw server with commands at once. It took a minute or so to recognize the error and correct for it. It required Andrew to stop the IDL script, and then Rick to reset the guide stage. To make sure everything was OK, Rick executed "obsmode", no problems, handed back to Andrew and Doug

0603 Finished taking data, over to Rick.

MOS Field Acquisition Test - Stone F Field

Just in time for the active optics experiments to be done, the wind is starting to pick up...

0608 - Wind gusting to 18m/s so we're moving the enclosure into a better position to gauge the true wind speed before proceeding.

0620 - Wind gusting over 20m/s limit, David is closing for now, telescope to Horizon-pointing

0830 - Winds still ~20m/s with gusts up to 23m/s in the past hour. Clouds rolled by briefly, then clear again.

1030 - Winds now average ~20m/s, with gusts above 25m/s rattling the building, so we're still closed and likely to stay closed for the rest of the night.

1130 - Winds still too high to open. We've put the instrument to sleep and called it a night.

-- RichardPogge - 17 Feb 2011
Topic revision: r9 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback